Monday May 23
The day shift began with the Antiproton Source (Pbar) stack at 103.5 E10, Tevatron (TeV) studies, and with Pbar stacking, and with MiniBooNE, NuMI, and Meson MCenter taking beam.
Shot setup begins at 10:18 AM.
Operations established store 4171 at 12:47 PM with an initial luminosity of 70.89E30. Pbar resumed stacking, beam to MCenter, and NuMI.
Operations put the Recycler Ring (RR) into shot setup at 7:11 PM to stash antiprotons.
After encountering many problems, the Recycler transfers were completed at 11:15 PM. Some of the problems encountered were a power supply (D:V901) tripping off, low level RF not matched between the Main Injector (MI) and the Recycler, and problems with the Recycler Ring closure.
Tuesday May 24
The midnight shift began with Operations monitoring store 4171 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.
Store 4171 was terminated at 2:00 PM.
Shot setup to the Tev begins at 2:19 PM.
Problems with a power supply (M:V205), and the failure of the F4 Wet engine fly wheel delayed shot setup.
Operations established store 4172 at 6:36 PM with an initial luminosity of 96.57E30. Pbar resumed stacking, beam to MCenter, and NuMI.
Store 4172 was lost at 10:56 PM, when all of the Tev RF stations tripped off. No cause was found for the trips.
Wednesday May 25
The midnight shift began with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.
Shot setup to the Tev begins at 3:12 AM.
Problems with the Main Injector Low level RF required a reboot of the RF system.
Problems with the frequency setting for the Pbar trombone also delayed shot setup.
The Plan for Wednesday
Establish a Tevatron store and begin stacking once again.
Accelerator Update Archive
For Tevatron luminosity charts and the current status of Fermilab's
accelerators and detectors (live!), please go to Fermilab Now
Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: firstname.lastname@example.org.